# Transition Metal Dichalcogenides Multijunction Solar Cells Toward the Multicolor Limit

### v1

@@ 3.5 Entropy loss as upward-emitted luminescence power and coupling thermalization
A recurring conceptual challenge is to interpret luminescence in multijunction stacks: luminescence is both a necessity for high $V_{\mathrm{oc}}$ and a potential loss channel if emitted into non-useful modes. To make this quantitative, we compute the upward-emitted luminescence power
$P^{\uparrow}_{\mathrm{lum}}=\sum_{i}\int E\,\phi^{\uparrow}_{\mathrm{em},i}(E,V_{i})\,dE,$ (14)
and report $P^{\uparrow}_{\mathrm{lum}}/P_{\mathrm{in}}$ as an entropy-loss proxy for emission-channel losses. In reciprocal stacks without intermediate mirrors, $P^{\uparrow}_{\mathrm{lum}}$ is nonzero because the stack must emit externally. In the idealized nonreciprocal model, $P^{\uparrow}_{\mathrm{lum}}$ can be suppressed in multijunction architectures, enabling additional headroom.13
Downward luminescence, in contrast, is not necessarily lost: it can be absorbed by the next junction and partially converted to electrical work. However, the portion of downward luminescent power that exceeds the electrical work extracted by the lower junction becomes heat (thermalization). We compute a coupling-thermalization proxy,
$P_{\mathrm{therm}}^{\mathrm{LC}}=\sum_{i=1}^{N-1}\left(P^{\downarrow}_{\mathrm{lum},i}-V_{i+1}J_{\mathrm{LC},i}\right),$ (15)
where $P^{\downarrow}_{\mathrm{lum},i}$ is the downward luminescent power emitted by junction $i$ and $J_{\mathrm{LC},i}$ is the corresponding coupling current absorbed by junction $i+1$. In the Results, we use these definitions to separate upward entropy loss from coupling-induced thermalization for a representative $N=5$ ladder.
@@ 3.6 Finite absorptance, excitonic TMD absorption, and thickness constraints
While the SQ limit assumes unity absorptance above $E_{g}$, real vdW absorbers are ultrathin. For TMDs, absorption is dominated by strong excitonic resonances and large absorption coefficients in the visible; nevertheless, tens-of-nanometers thickness is typically required to approach unity absorptance over broad spectral windows without strong light trapping.17, 4 We therefore supplement the step-absorber model with a thickness-dependent absorptance model
$A(E;t)=1-\exp[-\alpha(E)L_{\mathrm{eff}}(E)],\quad L_{\mathrm{eff}}(E)=F(E)\,t,$ (16)
@@
where $\alpha(E)$ is the absorption coefficient and $F(E)$ is an optional path-length enhancement factor. To capture TMD-like spectral shape, we adopt a minimal excitonic absorption coefficient model consisting of an Urbach tail, a continuum onset, and A/B exciton peaks (see Supporting Information). Because reciprocity requires the same $A(E;t)$ in Eq. (6), thickness directly affects both $J_{\mathrm{sc}}$ and $V_{\mathrm{oc}}$.
@@
In the ultrathin regime, geometric-optics light-trapping factors (e.g., $4n^{2}$) may be overly optimistic. At an elementary level, the path-length enhancement factor is simply $F(E)=L_{\mathrm{eff}}(E)/t$, i.e., the ratio between the average optical distance traveled by photons inside the absorber and the physical thickness $t$: $F(E)=1$ describes a single-pass planar film, while $F(E)>1$ captures multi-pass propagation and light trapping (back reflectors, resonant cavities, scattering/angle randomization, etc.). In thick absorbers with randomized ray directions, the broadband average is bounded by the Yablonovitch limit $F\leq 4n^{2}$, where $n$ is the refractive index.19
@@
When the thickness becomes comparable to the wavelength, however, light is better viewed as coupling through a finite set of radiative and guided modes, and the number of usable optical channels grows with thickness. Yu, Raman, and Fan derived wave-optics bounds on the maximum achievable enhancement in nanophotonic textures as a function of thickness and in-plane periodicity,19 and Miller emphasized the more general thickness–functionality connection for broadband optics.10 To capture this thickness-limited behavior without committing to a specific nanophotonic design, we introduce a normalized broadband enhancement function $g(t)\in[0,1]$ and define a thickness-dependent path-length enhancement factor
$F_{\mathrm{proxy}}(t)=1+\left(4n^{2}-1\right)g(t),$ (17)
@@
so that $F_{\mathrm{proxy}}=1$ in the no-light-trapping limit and $F_{\mathrm{proxy}}\to 4n^{2}$ only when sufficient thickness (and optical degrees of freedom) is available. In this work we use $n=4.5$ as a representative TMD value and take a simple saturating proxy $g(t)=1-\exp(-t/t_{0})$ with $t_{0}=200~\mathrm{nm}$ (see Supporting Information). When using this proxy in Eq. (16) we approximate $F(E)\!\approx\!F_{\mathrm{proxy}}(t)$, and we also discuss how wavelength-resolved bounds can be incorporated when a specific architecture is known (see Supporting Information).
@@ 4 Results
Efficiency versus junction number and the role of the bandgap window. Figure 2 summarizes the detailed-balance efficiency versus junction number $N$ for (i) unconstrained bandgaps and (ii) the conservative TMD bandgap window (1.0–2.1 eV), under both 1 sun and full concentration. Under full concentration, the unconstrained efficiency increases from $\sim 40\%$ (single junction) to $\sim 70.7\%$ for $N=5$ and $\sim 84.5\%$ by $N=50$ (approaching the reciprocal multicolor limit). Imposing the TMD window yields a pronounced large-$N$ plateau: $\sim 61.5\%$ for $N=5$ and $\sim 63.4\%$ by $N=50$. This plateau reflects an intrinsic materials constraint: once the accessible bandgap set no longer spans the entire solar spectrum, additional junctions provide diminishing returns.
Physically, the reciprocal multicolor limit is approached when the spectrum assigned to each junction becomes narrower as N increases, suppressing thermalization while maintaining a high radiative voltage under concentration. In the conservative TMD window, photons below 1.0 eV (near-IR) and above 2.1 eV (blue/UV) are irrevocably lost to transmission, so increasing N primarily repartitions the same limited spectral band. The rapid saturation by N on the order of five therefore implies that, within TMD-only bandgaps, further progress hinges more on luminescence quality (ERE), emission-channel control, and ultrathin optics than on pushing junction count. For example, under full concentration the TMD-window ladder rises from $\eta\approx 54.2\%$ at $N=2$ to $\eta\approx 61.5\%$ at $N=5$, whereas increasing junction count to $N=10$ yields only $\eta\approx 63.4\%$ (Table 1).
@@
Efficiency limit versus accessible bandgap window. To quantify how strongly the accessible bandgap window constrains large-$N$ performance, Figure 3 maps the efficiency limit (here shown for $N=20$ at full concentration) versus $(E_{g,\min},E_{g,\max})$. The conservative TMD point (1.0–2.1 eV) lies far from an unconstrained optimum, explaining the plateau in Figure 2. This map provides a compact way to evaluate how expanding the accessible bandgap range (e.g., by including other vdW semiconductors beyond conventional TMDs, or by alloying/strain) could raise the plateau.
The map also highlights an asymmetry between extending the low-gap and high-gap edges. Lowering Eg,min generally yields a larger gain because the solar photon flux is high in the near-IR, whereas raising Eg,max mainly improves voltage and reduces high-energy thermalization for a smaller portion of the spectrum. This suggests that identifying a stable narrow-gap bottom absorber in the 0.8–1.0 eV range can provide disproportionate leverage, even if the top gap remains near 2 eV. Physically, the lower cutoff $E_{g,\min}$ directly sets the fraction of near-infrared photons that are transmitted without conversion; within a fixed-width window, extending $E_{g,\min}$ downward therefore provides a larger efficiency headroom than modestly increasing $E_{g,\max}$.
@@
$E_{g}^{(N=5)}\approx(2.10,\,1.78,\,1.50,\,1.24,\,1.00)\ \mathrm{eV}\quad\text{(top $\to$ bottom)}.$ (18)
This ladder is used throughout subsequent analyses of luminescence, coupling, and thickness. Layer-resolved operating metrics for this representative $N=5$ stack (including ERE and finite-absorptance effects) are provided in the Supporting Information (Figures S1 and S2).
For the representative N=5 ladder, the rung spacing is not uniform. The optimizer concentrates rungs where the solar spectrum has high photon flux, while keeping a comparatively wide spacing between the lowest two bandgaps. This nonuniformity matters beyond spectral splitting: in reciprocal stacks without intermediate mirrors, smaller bandgap spacing increases downward luminescent coupling and the associated thermalization burden in the lower cell. Thus, optical design choices that control two-sided emission can be as important as the bandgap selection itself.
@@
Material mapping: candidate vdW/TMD absorbers for the $N=5$ ladder. To connect target bandgaps to realizable absorbers, Figure 5 overlays representative optical gaps reported for a selection of vdW semiconductors and TMD thickness variants.4, 17 The mapping is necessarily approximate because optical gaps depend on thickness, strain, dielectric environment, and exciton binding energy. Nevertheless, this comparison provides a practical starting point for selecting subcell candidates for transfer-printed stacks.
Two practical subtleties deserve emphasis when translating target optical gaps into device layers. First, monolayer TMD optical gaps are excitonic and depend on dielectric environment and strain, so the same nominal absorber can shift in both absorptance and radiative efficiency depending on encapsulation and adjacent layers. Second, many multilayer TMDs become indirect-gap as thickness increases, which can reduce luminescence and therefore reduce voltage unless light management and passivation mitigate nonradiative recombination.
Figure 5: Illustrative mapping between target bandgaps and vdW/TMD candidates. Symbols denote representative optical gaps drawn from the vdW/TMD literature (see Supporting Information, Table S3 for notes and sources), and dashed vertical lines denote the representative $N=5$ target ladder (Eq. (18)).
Table 1 provides a compact rung-to-material mapping for the representative $N=5$ ladder, emphasizing TMD-first choices within the conservative 1.0–2.1 eV window. The broader candidate database, including thickness/strain/alloy tunability ranges and optical-constant notes used in the thickness-dependent calculations, is provided in the Supporting Information (Tables S3–S6).
Among the five rungs, the bottom cell target near 1.0 eV is the most challenging within conventional TMDs and is also the most consequential for the efficiency limit because it controls near-IR photon capture. This makes contact engineering, encapsulation, and selection of narrow-gap vdW absorbers central to realizing the theoretical headroom.4, 17, 20
Table 1: Representative rung-to-material mapping for the $N=5$ target ladder within the conservative 1.0–2.1 eV window. Primary choices emphasize widely studied group-VI TMDs compatible with transfer printing; alternates illustrate thickness/strain/alloy routes and broader vdW options. See the Supporting Information, Tables S3–S6 for a more comprehensive candidate library and parameter notes.
@@
System-level penalty from finite ERE. Figure 6 shows how efficiency degrades as $\mathrm{ERE}$ decreases below unity, for selected $N$ values in the conservative TMD window. Because the voltage penalty is logarithmic, efficiencies remain relatively robust for $\mathrm{ERE}\gtrsim 10^{-2}$ but decline rapidly for lower $\mathrm{ERE}$. Importantly, in multijunction systems the penalty compounds across junctions: one low-ERE subcell can dominate the system loss.11, 14, 7 This motivates aggressive light-management and contact designs (low parasitic absorption, high-reflectivity mirrors, and low-loss transparent electrodes) that preserve external photon extraction in every subcell.
@@
A simple quantitative rule-of-thumb illustrates the compounding nature of the loss: at room temperature, an ERE of 1e-2 corresponds to an open-circuit voltage penalty on the order of one tenth of a volt per junction, so a five-junction stack can lose several tenths of a volt in total even if current matching is ideal. This is why maintaining high ERE in every subcell (not only the top) is a prerequisite for approaching multicolor efficiencies. For the representative $N=5$ ladder, reducing ERE from 1 to $10^{-2}$ lowers the full-concentration stack efficiency by about six absolute percentage points in the thick-limit regime (Figure 7).
@@
Figure 6: Efficiency versus external radiative efficiency (ERE) in the TMD window. Efficiency computed for $N=1,2,5$ under 1 sun (solid) and full concentration (dashed), illustrating the SQ-triangle voltage penalty and its correspondence to the thermodynamic $Q^{\mathrm{LED}}_{e}$ term.15, 14
@@
Thickness dependence with excitonic TMD absorptance. Figure 7 moves beyond step absorbers to incorporate finite absorptance in ultrathin vdW layers. Panel (a) shows the excitonic absorption coefficient model used here, including A/B exciton peaks and a continuum onset; the model is intentionally minimal but captures the qualitative spectral features of many group-VI TMDs.4, 17 The underlying optical-constant conventions ($n$, $k$, and $\alpha$) and representative material parameters used in this work are summarized in the Supporting Information (Tables S2–S6). Panel (b) shows the resulting $N=5$ stack efficiency versus a uniform subcell thickness $t$ for $\mathrm{ERE}=1$ and $\mathrm{ERE}=10^{-2}$. Even though monolayer and few-layer TMDs exhibit pronounced excitonic absorption near the band edge, the usable solar spectrum spans a broad range of above-gap energies and the continuum absorption remains finite; consequently, a very thin film does not absorb all incident photons in a single pass. In the Beer–Lambert regime ($\alpha t\!\ll\!1$), the absorptance scales approximately as $A\approx\alpha t$, so the short-circuit current and efficiency rise only gradually with thickness.
Using a 95% convergence criterion relative to the thick-absorber (step) limit for the full-concentration $N=5$ ladder, the required thickness is $\sim 120$ nm per subcell for a single-pass geometry ($F=1$). When broadband light trapping is included through the thickness-dependent proxy of Eq. (17)—i.e., $F(E)\approx F_{\mathrm{proxy}}(t)>1$ constructed from the normalized enhancement $g(t)$ in Figure 8c—this target thickness drops to $\sim 20$ nm. This shift provides an experimentally meaningful design message: moderate broadband path-length enhancement can trade optical thickness for physical thickness in ultrathin TMD stacks.
@@
Finite absorptance imposes a double penalty in detailed balance. First, incomplete absorption directly reduces $J_{\mathrm{sc}}$ by allowing above-gap photons to transmit. Second, because the same absorptance weights radiative emission channels, reduced absorption increases the effective recombination-to-generation ratio that sets the voltage (the $J_{\mathrm{sc}}/J_{0}$ ratio), lowering $V_{\mathrm{mpp}}$ and the overall power. The proxy $F_{\mathrm{proxy}}(t)$ accelerates convergence by increasing the effective photon dwell time, but it cannot make an arbitrarily thin absorber behave as a thick one: the normalized enhancement $g(t)$ is itself thickness-limited by nanophotonic channel counting and thickness–functionality bounds (Figure 8).19, 10, 6
@@
Nanophotonic bounds and the role of thickness. The need for a thickness-dependent broadband proxy follows from the distinction between geometric optics and wavelength-scale optics. In thick absorbers with randomized ray directions, the average internal path length can be increased up to the familiar Yablonovitch bound $F\leq 4n^{2}$. When $t\lesssim\lambda$, however, the optical response is mediated by a finite set of radiative and guided modes; the number of accessible channels (and therefore the achievable broadband response) grows with thickness. As a result, nanophotonic textures may yield strong enhancement at selected wavelengths or angles, but sustaining large enhancement over the broad solar spectrum is fundamentally thickness-limited.
@@
Figure 8 summarizes representative results from nanophotonic light-trapping theory and motivates the normalized enhancement function $g(t)$ used throughout this work. Panel (a) illustrates the Yu–Raman–Fan wave-optics enhancement oscillations around the $4n^{2}$ limit for periodic textures, highlighting the discrete-channel nature of wavelength-scale light trapping.19 Panel (b) illustrates single-mode scaling trends emphasized in nanophotonic bounds.19 Panel (c) plots the simple saturating function $g(t)$ adopted here, which can be interpreted as the fraction of the $4n^{2}$ enhancement that is experimentally achievable as a broadband average at thickness $t$. We then construct the effective path-length enhancement factor $F_{\mathrm{proxy}}(t)=1+(4n^{2}-1)g(t)$ [Eq. (17)], which is used in the absorptance model of Eq. (16) and in the thickness sweeps of Figure 7.10, 6
For multijunction stacks, this thickness constraint is wavelength selective: the longest-wavelength subcells are the hardest to enhance in the ultrathin regime, so the bottom junction often dictates the minimum practical thickness or the required strength of light trapping. Consequently, optical design should be co-optimized with the bandgap ladder, rather than treated as a post-processing step.10
@@
Figure 8: Nanophotonic bounds and a broadband path-length enhancement proxy. (a) Channel-counting proxy illustrating deviations from the geometric-optics $4n^{2}$ limit in the wavelength-scale regime. (b) Single-mode scaling proxy for $F_{\max}/(4n^{2})$ versus normalized thickness $\tau=d/(\lambda/2n)$. (c) Normalized broadband enhancement $g(t)=(F_{\mathrm{proxy}}(t)-1)/(4n^{2}-1)$ used to construct the thickness-dependent path-length enhancement factor $F_{\mathrm{proxy}}(t)=1+(4n^{2}-1)g(t)$ in Figure 7.19, 10, 6

### v3

@@ 3.5 Entropy loss as upward-emitted luminescence power and coupling thermalization
A recurring conceptual challenge is to interpret luminescence in multijunction stacks: luminescence is both a necessity for high $V_{\mathrm{oc}}$ or a potential loss channel (if it can be emitted into parasitic modes). To make this quantitative, we compute the upward-emitted luminescence power
$P^{\uparrow}_{\mathrm{lum}}=\sum_{i}\int E\,\phi^{\uparrow}_{\mathrm{em},i}(E,V_{i})\,dE$ (14)
and report $P^{\uparrow}_{\mathrm{lum}}/P_{\mathrm{in}}$ as an entropy-loss proxy for emission-channel losses. In reciprocal stacks without intermediate mirrors, $P^{\uparrow}_{\mathrm{lum}}$ is nonzero, because the stack must emit externally. In the idealized nonreciprocal model, $P^{\uparrow}_{\mathrm{lum}}$ can be suppressed in multijunction architectures, enabling additional efficiency gain.[11]
Downward luminescence, in contrast, is not necessarily lost: it can be absorbed by the next junction and partially converted to electrical work. However, the portion of downward luminescent power that exceeds the electrical work extracted by the lower junction becomes heat (i.e., thermalization loss). We compute a coupling-thermalization proxy:
$P_{\mathrm{therm}}^{\mathrm{LC}}=\sum_{i=1}^{N-1}\left(P^{\downarrow}_{\mathrm{lum},i}-V_{i+1}J_{\mathrm{LC},i}\right)$ (15)
where $P^{\downarrow}_{\mathrm{lum},i}$ is the downward luminescent power emitted by junction $i$, while $J_{\mathrm{LC},i}$ is the corresponding coupling current absorbed by junction $i+1$. We use these definitions to separate upward entropy loss from coupling-induced thermalization for a representative $N=5$ ladder.
@@ 3.6 Finite absorptance, excitonic TMD absorption, and thickness constraints
Although the SQ limit assumes unity absorptance above $E_{g}$, real TMD absorbers are ultrathin (less than 1 nm for a monolayer); consequently, their absorption cannot be unity. Also, it is dominated by strong excitonic resonances and large absorption coefficients in the visible; nevertheless, tens-of-nanometers thickness is typically required to approach unity absorptance over broad spectral windows without strong light trapping.[15, 3] We therefore supplement the step-absorber model with a thickness-dependent absorptance model:
$A(E;t)=1-\exp[-\alpha(E)L_{\mathrm{eff}}(E)],\quad L_{\mathrm{eff}}(E)=F(E)\,t$ (16)
@@
where $\alpha(E)$ and $F(E)$ are the absorption coefficient and an optional path-length enhancement factor, respectively. To capture TMD-like spectral shape, we adopt a minimal excitonic absorption coefficient model consisting of an Urbach tail, a continuum onset, and A/B exciton peaks (see Supplement Information). Because reciprocity requires the same $A(E;t)$ in Eq. (6), thickness directly affects both $J_{\mathrm{sc}}$ and $V_{\mathrm{oc}}$.
@@
In the ultrathin regime, geometric-optics light-trapping factors (e.g., $4n^{2}$, where $n$ is the refractive index) may be overly optimistic. At an elementary level, the path-length enhancement factor is simply $F(E)=L_{\mathrm{eff}}(E)/t$, i.e., the ratio between the average optical distance traveled by photons inside the absorber and the physical thickness $t$: $F(E)=1$ describes a single-pass planar film, while $F(E)>1$ reflects multi-pass propagation and light trapping (back reflectors, resonant cavities, and scattering/angle randomization). In thick absorbers with randomized ray directions, the broadband average is bounded by the Yablonovitch limit $F\leq 4n^{2}$.[17]
@@
When the thickness becomes comparable to the wavelength, however, light is better viewed as coupling through a finite set of radiative and guided modes, and the number of usable optical channels grows with thickness. Yu, Raman, and Fan derived wave-optics bounds on the maximum achievable enhancement in nanophotonic textures as a function of thickness and in-plane periodicity,[17] and Miller emphasized the more general thickness–functionality connection for broadband optics.[8] To capture this thickness-limited behavior without committing to a specific nanophotonic design, we introduce a normalized broadband enhancement function $g(t)\in[0,1]$ and define a thickness-dependent path-length enhancement factor:
$F_{\mathrm{proxy}}(t)=1+\left(4n^{2}-1\right)g(t)$ (17)
@@
so that $F_{\mathrm{proxy}}=1$ in the no-light-trapping limit and $F_{\mathrm{proxy}}\to 4n^{2}$ only when sufficient thickness (and optical degrees of freedom) is available. In this work, we use $n=4.5$ as a representative TMD $n$ value (in general, $n=4.5$ of representative TMDs are ranged from 4.17 to 4.70, as summarized in Table S2, Supplement Information) and take a simple saturating proxy $g(t)=1-\exp(-t/t_{0})$ with $t_{0}=200~\mathrm{nm}$ (see Supplement Information). When using this proxy in Eq. (16), we approximate $F(E)\!\approx\!F_{\mathrm{proxy}}(t)$, and also discuss how wavelength-resolved bounds can be incorporated, when a specific device architecture is known (see Supplement Information).
@@ 4 Results
Efficiency versus junction $N$ and the role of the bandgap window. Figure 2 summarizes the detailed-balance efficiency versus junction number $N$ for (i) unconstrained bandgaps and (ii) the conservative TMD bandgap window (1.0–2.1 eV), under both 1 sun and full concentration. Under full concentration, the unconstrained efficiency increases from $\sim 40\%$ (single junction) to $\sim 70.7\%$ for $N=5$ and $\sim 84.5\%$ by $N=50$ (approaching the reciprocal multicolor limit). Imposing the TMD window yields a pronounced large-$N$ plateau: $\sim 61.5\%$ for $N=5$ and $\sim 63.4\%$ by $N=50$. This plateau reflects an intrinsic materials constraint: once the accessible bandgap set no longer spans the entire solar spectrum, additional junctions provide diminishing returns.
Physically, the reciprocal multicolor limit is reached, when the spectrum assigned to each junction becomes narrower as $N$ increases. This multijunction approach suppresses thermalization, while maintaining a high radiative voltage under concentration. In the conservative TMD window, photons below 1.0 eV (near-IR) and above 2.1 eV (blue/UV) are irreversibly lost to transmission, so increasing $N$ primarily repartitions the same limited spectral band. The rapid saturation by $N$ on the order of five, therefore, indicates that, within TMD-only bandgaps, further progress hinges more on luminescence quality (ERE), emission-channel control, and ultrathin optics than on pushing junction count. For example, under full concentration, the TMD-window ladder rises from $\eta\approx 54.2\%$ at $N=2$ to $\eta\approx 61.5\%$ at $N=5$, whereas increasing junction $N$ to $10$ enables only $\eta\approx 63.4\%$ (Table 1).
@@
Efficiency limit versus accessible bandgap window. To quantify how strongly the accessible bandgap window constrains large-$N$ performance, Figure 3 maps the efficiency limit (herein, shown for $N=20$ at full concentration) versus $(E_{g,\min},E_{g,\max})$. The conservative TMD bandgap point (1.0–2.1 eV) are located far from an unconstrained optimum, explaining the plateau in Figure 2. This map provides a compact way to evaluate how expanding the accessible bandgap range (e.g., by including other vdW semiconductors beyond conventional TMDs, or by alloying/strain) could raise the efficiency plateau.
The map also highlights an asymmetry between extending the lowest- and highest-bandgap bounds. Lowering $E_{g,\min}$ generally leads to a larger efficiency gain, because the solar photon flux is relatively high in the near-infrared (NIR), whereas raising $E_{g,\max}$ mainly improves voltage and reduces high-energy thermalization. This suggests that identifying a stable narrow-gap bottom absorber in the 0.8–1.0 eV range can provide disproportionate leverage, even if the top bandgap remains near 2 eV. Physically, the lower cutoff $E_{g,\min}$ directly sets the fraction of NIR photons that are transmitted without conversion; within a fixed-width window, extending $E_{g,\min}$ downward, therefore, provides a larger efficiency gain potential than modestly increasing $E_{g,\max}$.
@@
$E_{g}^{(N=5)}\approx(2.10,\,1.78,\,1.50,\,1.24,\,and\ 1.00)\ \mathrm{eV}\quad\text{(from top to bottom)}$ (18)
This ladder is used throughout subsequent analyses of luminescence, coupling, and thickness. Layer-resolved operating metrics for this representative $N=5$ stack (including ERE and finite-absorptance effects) are provided in the Supplement Information (Figures S1 and S2).
For the representative $N=5$ ladder, the rung spacing is not uniform. The optimizer concentrates rungs, where the solar spectrum has relatively high photon flux, while keeping a comparatively wide spacing between the lowest two bandgaps. This nonuniformity matters beyond spectral splitting: in reciprocal stacks without intermediate mirrors, smaller bandgap spacing increases downward luminescent coupling and the associated thermalization burden in the lower cell. Thus, optical design choices that control two-sided emission can be as important as the bandgap selection itself.
@@
Material mapping: candidate vdW and TMD absorbers for the $N=5$ ladder. To connect target bandgaps to realizable absorbers, Figure 5 overlays representative optical bandgaps, which were reported for a selection of vdW semiconductors and TMD thickness variants.[3, 15] This mapping is necessarily approximate, because optical bandgaps depend on thickness, strain, dielectric environment, and exciton binding energy. Nevertheless, this comparison still provides a practical starting point for selecting TMD subcell candidates for stacks.
Two practical subtleties are noteworthy, when translating target optical bandgaps into device layers. First, monolayer TMD optical bandgaps are excitonic and depend on dielectric environment and strain, so the same nominal absorber can shift in both absorptance and radiative efficiency, depending on encapsulation and adjacent layers. Second, many multilayer TMDs tend to retain indirect-bandgap as thickness increases, which can reduce luminescence and therefore reduce voltage, unless light management and passivation mitigate nonradiative recombination. Among the five rungs, the bottom cell target near 1.0 eV is the most challenging within conventional TMDs and is also the most consequential for the efficiency limit, because it controls NIR photon harvesting. This makes contact engineering, encapsulation, and selection of narrow-gap vdW absorbers central to realizing the theoretical efficiency gain.[3, 15, 18] Table 1 provides a compact rung-to-material mapping for the representative $N=5$ ladder, emphasizing TMD-first choices within the conservative 1.0–2.1 eV window. The broader candidate database, used in the thickness-dependent calculations, such as thickness, strain, alloy tunability ranges, and optical-constants, is provided in the Supplement Information (Tables S3–S6).
Figure 5: Illustrative mapping between target bandgaps and vdW/TMD candidates. Symbols denote representative optical gaps drawn from the vdW/TMD literature (see Supplement Information, Table S3 for notes and sources), and dashed vertical lines denote the representative $N=5$ target ladder (Eq. (18)).
Table 1: Representative rung-to-material mapping for the $N=5$ target ladder within the conservative 1.0–2.1 eV window. Primary choices emphasize widely studied group-VI TMDs compatible with transfer printing; alternates illustrate thickness/strain/alloy routes and broader vdW options. See the Supplement Information, Tables S3–S6 for a more comprehensive candidate library and parameter notes.
@@
System-level penalty from finite ERE. Figure 6 shows how efficiency degrades as $\mathrm{ERE}$ decreases below unity, for selected $N$ values in the conservative TMD window. Because the voltage penalty is logarithmic, efficiencies remain relatively robust for $\mathrm{ERE}\gtrsim 10^{-2}$, but decline rapidly for lower $\mathrm{ERE}$. Importantly, in multijunction systems, the penalty compounds across junctions: one low-ERE subcell can dominate the system loss.[9, 12, 6] This motivates aggressive light-management and contact designs (e.g., low parasitic absorption, high-reflectivity mirrors, and low-loss transparent electrodes) that preserve external photon extraction in every TMD subcell.
@@
A simple quantitative rule-of-thumb illustrates the compounding nature of the loss: at room temperature, an ERE of 1e-2 corresponds to an open-circuit voltage penalty on the order of one tenth of a volt per junction, so a five-junction stack can lose several tenths of a volt in total, even if current matching is ideal. This is why maintaining high ERE in every TMD subcell (not only the top) is a prerequisite for approaching multicolor efficiencies. For the representative $N=5$ ladder, reducing ERE from 1 to $10^{-2}$ lowers the full-concentration stack efficiency by about six absolute percentage points in the thick-limit regime (Figure 7).
@@
Figure 6: Efficiency versus external radiative efficiency (ERE) in the TMD window. Efficiency computed for $N=1,2,5$ under 1 sun (solid) and full concentration (dashed), illustrating the SQ-triangle voltage penalty and its correspondence to the thermodynamic $Q^{\mathrm{LED}}_{e}$ term.[13, 12]
@@
Thickness dependence with excitonic TMD absorptance. Figure 7 moves beyond step absorbers to incorporate finite absorptance in ultrathin TMD layers. Panel (a) shows the excitonic absorption coefficient model used herein, including A/B exciton peaks and a continuum onset; the model is intentionally minimal but captures the qualitative spectral features of many group-VI TMDs.[3, 15] The underlying optical-constant conventions ($n$, $k$, and $\alpha$) and representative material parameters used in this work are summarized in the Supplement Information (Tables S2–S6). Panel (b) shows the resulting $N=5$ stack efficiency versus a uniform TMD subcell thickness $t$ for $\mathrm{ERE}=1$ and $\mathrm{ERE}=10^{-2}$. Even though monolayer and few-layer TMDs exhibit pronounced excitonic absorption near the band edge, the usable solar spectrum spans a broad range of above-bandgap energies and the continuum absorption remains finite; consequently, a very thin film does not absorb all incident photons in a single pass. In the Beer–Lambert regime ($\alpha t\!\ll\!1$), the absorptance scales approximately as $A\approx\alpha t$, so the short-circuit current and efficiency rise only gradually with thickness.
@@
Using a 95% convergence criterion relative to the thick-absorber (step) limit for the full-concentration $N=5$ ladder, the required thickness is $\sim 120$ nm per TMD subcell for a single-pass geometry ($F=1$). When broadband light trapping is included through the thickness-dependent proxy of Eq. (17) (i.e., $F(E)\approx F_{\mathrm{proxy}}(t)>1$ constructed from the normalized enhancement $g(t)$ in Figure 8c), this target thickness drops to $\sim 20$ nm. This shift provides an experimentally important design message: moderate broadband path-length enhancement can trade ”optical” thickness for ”physical” thickness in ultrathin TMD stacks. Finite absorptance imposes a double penalty in detailed balance. First, incomplete absorption directly reduces $J_{\mathrm{sc}}$ by allowing above-bandgap photons to transmit. Second, because the same absorptance weights radiative emission channels, reduced absorption increases the effective recombination-to-generation ratio that sets the voltage (the $J_{\mathrm{sc}}/J_{0}$ ratio) and lowers both $V_{\mathrm{mpp}}$ and the overall power. The proxy $F_{\mathrm{proxy}}(t)$ accelerates convergence by increasing the effective photon dwell time, but it cannot make an arbitrarily thin absorber behave as a thick one: the normalized enhancement $g(t)$ is itself thickness-limited by nanophotonic channel counting and thickness–functionality bounds (Figure 8).[17, 8, 5]
@@
Nanophotonic bounds and the role of thickness. The need for a thickness-dependent broadband proxy follows from the distinction between geometric optics and wavelength-scale optics. In thick absorbers with randomized ray directions, the average internal path length can be increased up to the Yablonovitch bound $F\leq 4n^{2}$. When $t\lesssim\lambda$, however, the optical response is mediated by a finite set of radiative and guided modes; the number of accessible channels (and therefore the achievable broadband response) grows with thickness. As a result, nanophotonic textures may yield strong enhancement at selected wavelengths or angles, but sustaining large enhancement over the broad solar spectrum is fundamentally thickness-limited.
@@
Figure 8 summarizes representative results from nanophotonic light-trapping theory and motivates the normalized enhancement function $g(t)$ used throughout this work. Figure 8(a) illustrates the Yu–Raman–Fan wave-optics enhancement oscillations around the $4n^{2}$ limit for periodic textures, highlighting the discrete-channel nature of wavelength-scale light trapping.[17] Figure 8(b) illustrates single-mode scaling trends emphasized in nanophotonic bounds.[17] Figure 8(c) plots the simple saturating function $g(t)$ adopted here, which can be interpreted as the fraction of the $4n^{2}$ enhancement that is experimentally achievable as a broadband average at thickness $t$. We then construct the effective path-length enhancement factor $F_{\mathrm{proxy}}(t)=1+(4n^{2}-1)g(t)$ [Eq. (17)], which is used in the absorptance model of Eq. (16) and in the thickness sweeps of Figure 7.[8, 5] For multijunction stacks, this thickness constraint is wavelength selective: the longest-wavelength TMD subcells are the hardest to enhance in the ultrathin regime, so the bottom junction often dictates the minimum practical thickness or the required strength of light trapping. As such, optical design should be co-optimized with the bandgap ladder, rather than treated as a post-processing step.[8]
@@
Figure 8: Nanophotonic bounds and a broadband path-length enhancement proxy. (a) Channel-counting proxy illustrating deviations from the geometric-optics $4n^{2}$ limit in the wavelength-scale regime. (b) Single-mode scaling proxy for $F_{\max}/(4n^{2})$ versus normalized thickness $\tau=d/(\lambda/2n)$. (c) Normalized broadband enhancement $g(t)=(F_{\mathrm{proxy}}(t)-1)/(4n^{2}-1)$ used to construct the thickness-dependent path-length enhancement factor $F_{\mathrm{proxy}}(t)=1+(4n^{2}-1)g(t)$ in Figure 7.[17, 8, 5]